OpenStudy (anonymous):

can someone ACTUALLY help me with 7th grade algebra?

OpenStudy (anonymous):

what is your question

OpenStudy (anonymous):

ikr?

OpenStudy (anonymous):

well, everything. i can't do it and hey someone >:(

OpenStudy (anonymous):

somebody*

OpenStudy (anonymous):

i need help with variables?

OpenStudy (anonymous):

ok we're ready

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i don't have anything, i just need help. i don't have an exuasion ready

OpenStudy (anonymous):

equation*

OpenStudy (anonymous):

5x=20 /5 /5 x=4\=

OpenStudy (anonymous):

2x-14=10 +14 +14 2x=24 /2 /2 x=12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

have a good day

OpenStudy (anonymous):

hey sara u ther?